Given four vector s at equilibrium: T, N, F, and W in a 3 dimensional space Point O is at the origin of the coordinate system and has coordinates of (0, 0,0) Given that T is parallel to OB, N isparallel to OA, Fis parallel to OC, and W is a given vector of coordinate 〈 0,-5,0 〉 use unit vectors to calculate the magnitude of T, F, and N

A (2,5,-1) B-6,3,4) c (2,-1-4)
Answer #1

magnitude = sqrt_{x^2+y^2+z^2}

T = |T| x OB / |OB|

= T (-6i + 3j + 4k) / V6+32+42

= -0.77Ti + 0.385Tj + 0.51Tk

similarly F = 0.44Fi - 0.22Fj -0.88Fk

and N = 0.36Ni + 0.9Nj - 0.18Nk

W = -5j

As the system is in equilibrium

sum of all vectors should be zero. Hence

T+F+N+W = 0

-0.77T + 0.44F +0.36N = 0

0.385T - 0.22F + 0.9N = 5

0.51T -0.88F -0.18N = 0

solving above 3 eqns

T = 2.43

F = 0.46

N = 4.63